Question 7 A heat lamp emits infrared radiation whose rms electric field is E 3500 N/C. (a) What is the average intensity of the radiation? (b) The radiation is focused on a persons leg over a circular area of radius 3.3 cm. What is the average power delivered to the leg? (c) The portion of the leg being irradiated has a mass of 0.20 kg and a specific heat capacity of 3500 J/(kg C How long does it take to raise its temperature by 1.5 C°? Assume that there is no other heat transfer into or out of the portion of the leg being heated. Number
Answer #1

a) we know,

Intensity of radiation,
I = Erms^2/(c*mue)

= 3500^2/(3*10^8*4*pi*10^-7)

= 3.25*10^4 W/m^2

b) we know, Intensity = Power/Area

==> Power = Intensity*Area

= 3.25*10^4*pi*0.033^2

= 111.2 W

c) let t is the time taken.

Q = m*c*dT

P*t = m*c*dT

t = m*c*dT/P

= 0.2*3500*1.5/111.2

= 9.44 s